OpenStudy (anonymous):

Mia’s mother gives her $10 per week for allowance. Mia gets an additional $3 for every hour she spends watching her little brother. Mia wants to make at least $25 this week to spend shopping with her friends. Use this information to help you answer questions 6-8 below. Please explain your answers with words, and show your work, including computations. 1. Write an inequality, using variables, to represent he number of times that Mia has to watch her brother in order to get at least $25 2. Solve your inequality and describe the solution. 3. Create a visual model to represent

OpenStudy (anonymous):

helppppp please!!

OpenStudy (anonymous):

1. 10 + 3 x >= 25 2. 3x >= 15 x >= 5 she must watch her brother for at least 5 hours 3.12(2x + 5) – 3(-2x + 4) = 24x +60 +6x -12 = 30x +48

OpenStudy (anonymous):

Welcome!!!